# Non-local planelike minimizers and $\Gamma$-convergence of periodic energies to a local anisotropic perimeter

### v1

@@ Appendix C Behavior of $\operatorname{P}_{K}$ in large domains
Moreover, for almost every $x,y\in C_{\rho}$, we have that
$\begin{split}&|\widetilde{x}-\widetilde{y}|=|x-\psi(x)\phi(x^{\prime})e_{n}-y+\psi(y)\phi(y^{\prime})e_{n}|\\ &\qquad\leqslant|x-y|+|\psi(x)|\,|\phi(x^{\prime})-\phi(y^{\prime})|+|\phi(y^{\prime})|\,|\psi(x)-\psi(y)|\\ &\qquad\leqslant|x-y|+[\phi]_{Lip(B_{\rho}^{n-1})}|x^{\prime}-y^{\prime}|+[\phi]_{Lip(B_{\rho}^{n-1})}\rho\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}|x-y|\\ &\qquad\leqslant c_{2}|x-y|\end{split}$
for some constant $c_{2}>0$ independent of $\rho$. This yields
(C.9) $|x-y|^{-n-2s}\leqslant c_{2}|\widetilde{x}-\widetilde{y}|^{-n-2s},$
up to renaming $c_{2}$.
Using the change of variable $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$ together with (C.8), (C.7), and (C.9), we conclude that
$\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dxdy}{|x-y|^{n+2s}}\leqslant\frac{c_{2}}{1-c_{1}}\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{d\widetilde{x}d\widetilde{y}}{|\widetilde{x}-\widetilde{y}|^{n+2s}},$
showing (C.3). ∎
@@
Since $\Omega$ is bounded, there exists a constant $\zeta>0$ such that $\Omega\subseteq B_{\zeta}$. Up to scaling, we can suppose that $\zeta=1$, otherwise we consider $\widetilde{\eta}:=\eta\zeta\in(0,1)$.
Moreover, let $r_{\Omega}$, $L_{\Omega}$, and $\alpha_{\Omega}$ be as in Definition C.1. We define
@@
$3\sqrt{n}\delta<Rr_{\Omega}\quad\mbox{and}\quad 3\sqrt{n}\delta\lambda<R\alpha_{\Omega},$
@@
$(\Omega_{R}\times\Omega_{R}^{c})\cap\{|x-y|<\delta\}\subseteq\bigcup_{i,j=0}^{N_{\delta}}B_{\sqrt{n}\delta}(x_{i})\times B_{\sqrt{n}\delta}(x_{j})\cap\{|x-y|<\delta\}$
$\qquad\qquad\subseteq\bigcup_{i=0}^{N_{\delta}}B_{3\sqrt{n}\delta}(x_{i})\times B_{3\sqrt{n}\delta}(x_{i}).$
@@
$\partial(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i}))\cap C_{3\sqrt{n}\delta,\lambda}=\left\{(x^{\prime},\phi_{i}(x^{\prime}))\mbox{ s.t. }x^{\prime}\in B_{3\sqrt{n}\delta}^{n-1}\right\}.$
@@
(C.11) $\begin{split}\mathcal{I}_{1}&=\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|<\delta\}\end{subarray}}K(x,y)\,dx\,dy\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\Omega_{R}\cap B_{3\sqrt{n}\delta}(x_{i}))\times(\Omega_{R}^{c}\cap B_{3\sqrt{n}\delta}(x_{i}))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{3\sqrt{n}\delta})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.\end{split}$
@@
$\Phi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\delta,\lambda}$
and $\Phi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\delta,\lambda}.$
@@
(C.12) $\begin{split}&\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\qquad\qquad\leqslant c_{1}^{i}\iint_{(\{x_{n}<0\}\cap C_{3\sqrt{n}\delta})\times(\{x_{n}>0\}\cap C_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}},\end{split}$
for some positive constant $c_{1}^{i}$ depending only on $n$, $s_{1}$, and $\Phi_{i}$.
Moreover, recalling (C.6) and (C.9) from the proof of Lemma C.4, we infer that
$c_{1}^{i}\leqslant\left(1+[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}+\frac{2[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}{\lambda-[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}\right)^{-(n+2s_{1})}\frac{1}{1-\frac{9}{16}}\leqslant\frac{16}{7},$
for every $i$. Let us set
@@
$\Psi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}(x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(-3\sqrt{n}\eta\lambda R,0)$
and $\Psi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(0,3\sqrt{n}\eta\lambda R).$
@@
Then, using again Lemma C.4 and setting $c_{2}:=\max_{i}c_{2}^{i}$ (which does not depend on $R$), we have
(C.17) $\begin{split}&\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2}^{i}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{(|x^{\prime}-y^{\prime}|^{2}+|x_{n}-y_{n}|^{2})^{\frac{n+2s_{2}}{2}}},\end{split}$
@@
$\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant c_{2}R^{n-1}.$

### v2

@@ Appendix C Behavior of $\operatorname{P}_{K}$ in large domains
Moreover, for almost every $x$, $y\in C_{\rho}$, using again (C.3) and (C.4) we have that
(C.10) $\begin{split}|\widetilde{x}-\widetilde{y}|&=|x-\psi(x)\phi(x^{\prime})e_{n}-y+\psi(y)\phi(y^{\prime})e_{n}|\\ &\leqslant|x-y|+|\psi(x)|\,|\phi(x^{\prime})-\phi(y^{\prime})|+|\phi(y^{\prime})|\,|\psi(x)-\psi(y)|\\ &\leqslant|x-y|+[\phi]_{Lip(B_{\rho}^{n-1})}|x^{\prime}-y^{\prime}|+[\phi]_{Lip(B_{\rho}^{n-1})}\rho\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}|x-y|\\ &\leqslant c_{2}|x-y|\end{split}$
for some constant $c_{2}>0$ depending on $\lambda$ but independent of $\rho$.
Employing the changes of variable $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$ together with (C.8), (C.9), and (C.10), we conclude that
$\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dx\,dy}{|x-y|^{n+2s}}\leqslant\frac{c_{3}}{1-c_{1}}\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{d\widetilde{x}\,d\widetilde{y}}{|\widetilde{x}-\widetilde{y}|^{n+2s}},$
for some $c_{3}>0$ independent of $\rho$, showing the desired result. ∎
@@
Since $\Omega$ is bounded, there exists $\zeta>0$ such that $\Omega\subseteq B_{\zeta}$. Up to scaling, we can suppose that $\zeta=1$.
Moreover, let $r_{\Omega}$, $L_{\Omega}$, and $\alpha_{\Omega}$ be as in Definition C.1 and define
@@
$3\sqrt{n}\delta<Rr_{\Omega}\quad\mbox{and}\quad 3\sqrt{n}\delta\lambda<R\alpha_{\Omega}$
@@
$(\Omega_{R}\times\Omega_{R}^{c})\cap\{|x-y|<\delta\}$ $\subseteq\bigcup_{i,j=0}^{N_{\delta}}\Big((\Omega_{R}\cap B_{\sqrt{n}\delta}(x_{i}))\times(\Omega_{R}^{c}\cap B_{\sqrt{n}\delta}(x_{j}))\Big)\cap\{|x-y|<\delta\}$
$\subseteq\bigcup_{i=0}^{N_{\delta}}\Big((\Omega_{R}\cap B_{3\sqrt{n}\delta}(x_{i}))\times(\Omega_{R}^{c}\cap B_{3\sqrt{n}\delta}(x_{i}))\Big).$
@@
$\partial(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i}))\cap C_{3\sqrt{n}\delta,\lambda}=\left\{(x^{\prime},\phi_{i}(x^{\prime}))\mbox{ s.t. }x^{\prime}\in B_{3\sqrt{n}\delta}^{n-1}\right\}.$
@@
(C.12) $\begin{split}\mathcal{I}_{1}&=\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|<\delta\}\end{subarray}}K(x,y)\,dx\,dy\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\Omega_{R}\cap B_{3\sqrt{n}\delta}(x_{i}))\times(\Omega_{R}^{c}\cap B_{3\sqrt{n}\delta}(x_{i}))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap B_{3\sqrt{n}\delta})\times(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap B_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.\end{split}$
@@
$\Phi_{i}(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\delta,\lambda}$
and $\Phi_{i}(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\delta,\lambda}.$
@@
(C.13) $\begin{split}&\iint_{(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\qquad\qquad\leqslant c_{1,i}\iint_{(\{x_{n}<0\}\cap C_{3\sqrt{n}\delta})\times(\{x_{n}>0\}\cap C_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}},\end{split}$
for some positive constant $c_{1,i}$ depending only on $n$, $s_{1}$, and $\Phi_{i}$.
Moreover, recalling (C.7) and (C.10) from the proof of Lemma C.4, we infer that, for every $i$,
$c_{1,i}\leqslant\left(1+[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}+\frac{2[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}{\lambda-[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}\right)^{-(n+2s_{1})}\frac{1}{1-\frac{9}{16}}\leqslant\frac{16}{7}.$
Let us set
@@
$\Psi_{i}(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(-3\sqrt{n}\eta\lambda R,0)$
and $\Psi_{i}(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(0,3\sqrt{n}\eta\lambda R).$
@@
Then, using again Lemma C.4 and setting $c_{2}:=\max_{i}c_{2,i}$ (which does not depend on $R$), we have
(C.18) $\begin{split}&\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2,i}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{(|x^{\prime}-y^{\prime}|^{2}+|x_{n}-y_{n}|^{2})^{\frac{n+2s_{2}}{2}}},\end{split}$
@@
$\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}(\Omega_{R}-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}(\Omega_{R}^{c}-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant c_{2}R^{n-1}.$